Solve using the multiplication principle first. Then use the elimination method:
7p+5q=2 8p – 9q =17
can you use a high number to multiply like 45 to cancel out the q's help!

Answer by Alan3354(69443)

Solve using the multiplication principle first. Then use the elimination method:
7p + 5q=2
8p – 9q =17
can you use a high number to multiply like 45 to cancel out the q's help!
------------
Is 45 a "high number?" What about 5,678,932 ?
-------------
7p + 5q=2 Multiply by 9
8p – 9q =17 Multiply by 5
------------
63p + 45q = 18
40p - 45q = 85
---------------- Add
103p = 103
p = 1
--------
You can finish it.